### v2

# Faculty of Computer Science, TU Dresden, Germany

Abstract
We introduce the Riskman ontology & shapes for representing and analysing information about risk management for medical devices.
Risk management is concerned with taking necessary precautions so a medical device does not cause harms for users or the environment.
To date, risk management documentation is submitted to notified bodies (for certification) in the form of semi-structured natural language text.
We propose to use classes from the Riskman ontology to logically model risk management documentation, and to use the included SHACL constraints to check for syntactic completeness and conformity to relevant standards.
In particular, the ontology is modelled after ISO 14971 and the recently published VDE Spec 90025.
Our proposed methodology has the potential to save many person-hours for both manufacturers (when creating risk management documentation) as well as notified bodies (when assessing submitted applications for certification), and thus offers considerable benefits for healthcare and, by extension, society as a whole.
Resource type: OWL EL Ontology, SHACL Shapes
URL: https://w3id.org/riskman/
Keywords:
Risk management Safety assurance OWL EL SHACL1 Introduction
Medical devices typically are safety-critical, meaning their failure under certain conditions can lead to harm to humans or the environment. To ensure that potential harms are minimised, legislation requires manufacturers of medical devices to provide a comprehensive justification that their product is acceptably safe. In Europe, such legal requirements mainly stem from the European Union’s Medical Device Regulation (EU MDR) [1]. After its introduction in 2017, the original plan was that manufacturers have time until May 2024 to re-certify their devices under the new regulations. This caused an immense backlog for notified bodies (organisations certifying medical devices in Europe, e.g., TÜVs in Germany), so the EU decided to extend the transition period until at least 2027 to avert a potential shortage of medical devices (and ramifications thereof) [44].
The fundamental challenges of medical device certification are (1) the sheer amount of information that notified bodies have to process (even for a single device), and (2) the way this information is submitted to notified bodies, namely in the form of text. In principle, device manufacturers have to argue that they have proactively, systematically, and thoroughly analysed and mitigated the risks associated with their device as much as possible. One way to do this is using a so-called assurance case, a structured argument supported by a body of evidence that together provide a compelling, comprehensible, and valid case that a system is acceptably safe for a given application in a given context. Using assurance cases has a long(er) tradition in other areas of safety-critical systems, e.g. in aviation or nuclear power plants [67], but the need for assurance cases for medical devices has been recognised long ago [73]. The Food and Drug Administration of the U.S. (FDA) has even provided a guidance document for one class of products (infusion pumps) due to an unusually large number of previous incidents [69].
As much as the introduction of assurance cases has helped advance safety management practices in the medical device industry, there is still a lot to be wanted. Sujan et al. even claim that “safety management practices in healthcare [are] at present […] less mature than those in other safety-critical industries” [67, p. 185]. We briefly note some of the shortcomings we address in this work.
Since the intended recipient of a safety case is a human (auditor), a lot of work in safety engineering has focused on the presentation of the argument, e.g. to avoid confirmation bias [17]. In particular, several approaches to visualise assurance cases exist, e.g. the goal structuring notation [46]. There is much less work on the representation of safety arguments, since applicable legislation typically requires a submission in the form of text.
Due to the sheer volume of text submitted to notified bodies for certification, an auditor can never look at all relevant points in detail; furthermore, auditors typically spend a considerable proportion of their time navigating through documents (best case, using text search in a document viewer). Significant time savings are offered by semi-automatic checks and improved navigation through e.g. semantics-enhanced document search.
Reusing parts of safety arguments is a widely recognised problem [52, 60]. The simplest instance of re-use happens when a manufacturer continues development of a certified device and wants to certify the “updated version”: application lifecycle management tools can help identify the delta between two device versions and risk managers only need to concentrate on how that difference affects risks and mitigations. Harder instances of re-use are equally pervasive and important, e.g. when a manufacturer wants to use a part of a device in a novel device of a different type, or when a manufacturer buys parts (like integrated circuits) from component suppliers and needs to use assurances about the parts in order to claim overall device safety. We even propose to go beyond this and technically enable re-use across manufacturers, which is especially relevant for small/medium enterprises that may lack staff or experience in risk management.
Thus there is substantial demand for intelligent software support in risk management, both on the side of manufacturers as well as from notified bodies.
In this paper, we propose to use logical modelling to represent risk management documentation, and the shapes constraint language (SHACL) to check those representations for conformance with a set of requirements, e.g., whether all identified risks have an associated mitigation. While such straightforward checks could clearly also be achieved by custom software, the presence of a logical inference step before the constraint checking step is a significant advantage of our approach, as we shall demonstrate later in the paper. Taken together, the Riskman ontology and its shapes constraints constitute the resource we contribute in this paper, which is intended to be Riskman’s canonical citation. As far as we are aware, using OWL & SHACL is a novel approach to representing and reasoning about risk management documentation. The requirements to be checked against are simple and syntactic at the moment, but from communication with domain experts from notified bodies and manufacturers, we still expect our contributed resource and methodology to be a major step forward as manufacturers (and notified bodies) can expect to “save” the first round of conformity assessment, which can translate into a several-months-long reduction in time to market. On the other hand, our resource does not aim at assessing the adequacy and correctness of the implemented mitigations (such as whether an insulation thickness of 0.5mm is adequate to lower the risk of electric shock), which is presumably AI-complete and would require vast amounts of background knowledge. Apart from this important application domain, our ontology and shapes are also of independent interest to the Semantic Web community, as the combination of logical inference and SHACL constraint checking is a topic that has garnered research interest lately [4, 57, 22]. Finally, the use of our approach in the medical device industry and on the part of notified bodies entails a considerable impact on the adoption of Semantic Web technologies in general. By employing these technologies ourselves, we furthermore enable Riskman users to benefit from the good infrastructure of tooling and data the Semantic Web field provides. Risk reports can easily be queried or enhanced with additional information about the use context of a concrete device or data about particular patient problems by simply using other ontologies and data [55, 25, 54].
The rest of the paper is organised as follows: In the next section, we give an overview on related work, mostly ontologies involving notions of risk, but also software-based solutions for safety case management. Afterwards (Sect. 3), we introduce the state of the art of risk management and the basic notions introduced by the norms upon which we build our work, both in general terms and with an illustrative running example. Sect. 4 then introduces the Riskman ontology and its associated shape constraints, explaining how we intend them to be used in risk management, and showcasing them with the running example. In Sect. 5 we conclude with a discussion of potential future work.
2 Related Work
2.1 Conceptual Work and Ontologies
Fenz et al. [27, 26] created an ontology for the information security standard ISO 27002. While they do not deal with risk management (or medical devices), they also use their ontology along with a reasoner to infer information about compliance with the standard’s requirements. Reasoning results are then interpreted by humans or by a tailor-made software tool.
Uciteli et al. [68] provided the Risk Identification Ontology, which defines notions of risk and adverse situation and embeds them into the top-level ontology GFO (General Formal Ontology) [37, 38]. However, they work in a more process-oriented setting with special focus on risk identification in time periods surrounding surgical procedures. In that setting, risks cannot be mitigated beforehand, and so risk management – our main focus – is not within their scope.
Kim et al. [47] presented a process integration ontology for medical software developers with a focus on medical devices, combining notions from IEC 60601-1, IEC 62304, and, notably, ISO 14971. The resulting integrated ontology is however mainly designed to help developers comply with the involved standards; the ontology was not developed with logical reasoning as explicit intended use case.
Aziz et al. [9] developed a Hazard Identification Ontology, involving notions of Hazard and Events. Their focus was however on identifying risks rather than mitigating them, especially in scenarios of fire, explosions, or toxicity.
Schütz et al. [62] created an ontology for medical devices in Germany, albeit more broadly targeting devices’ manufacturers, operators, and legal procedures from an outside perspective with the aim of general semantic interoperability, and based on a legal framework that has since been superseded by the EU MDR.
Single et al. [64] presented an ontology for Hazard and Operability (HAZOP), a methodology for scenario-based hazard evaluation, therein defining notions of deviation, cause, effect, consequence, and safeguard. Their aim was however to create HAZOP worksheets (to be used by human operators) automatically.
Alanen et al. [5] provided a comprehensive risk assessment ontology including the notions risk, risk level, and risk control, which is harmonised between (and intended for use across) safety, security, and dependability. The intended use of their ontology is to “support the creation of a structured work product storage with traceability links” in order to improve upon current practices with non-structured word-processing documents, albeit without using logical reasoning. Conclusion: While various ontologies including notions of risk exist, they either do not implement ISO 14971 [40] or are not designed with logical reasoning in mind, and furthermore none of them uses SHACL for conformance testing.
2.2 (Logic-Based) Software Tools for Risk Management
Fujita et al. [29] presented the D-Case Editor, a dependability-oriented assurance case editor supporting the goal structuring notation [46], implemented as an Eclipse plugin and available at its GitHub repository [20]. However, its development has been discontinued (the last repository change was in 2015).
Rushby [61] and Cruanes et al. [19] introduced the Evidential Tool Bus, a Datalog-based system for integrating the development of safety cases into system (and software) development. In particular, their approach covers the management of claims (about the system to be developed) and how they are supported by evidence provided by other software tools. While an implementation is available [56], it is not maintained any longer (last change 2016). Beyene and Ruess [13] later picked up that work and extended it by connecting to the Jenkins [42] continuous integration software, with an active repository [12]. The focus of the Evidential Tool Bus is primarily on obtaining, maintaining, and managing pieces of evidence in a development setting. As such, it can be an important and complementary addition to using the methodology we propose in this paper.
Since the issue of safety is of great concern to industry, there are also several proprietary software tools centring around safety assurance: The UK company Adelard offers the Assurance and Safety Case Environment. Similarly, de la Vara et al. [21] reported on using the V&V studio by REUSE Software.
Conclusion: Existing software solutions for risk management are proprietary or do not use formal logics, or they have other (but complementary) use cases.
3 Background
| Term | Definition |
| Analysed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Controlled risk | Structured artifact that relates one analysed risk with one or more SDA(s) and specifies a resulting residual risk. |
| Domain specific hazard | Structured artifact that centres around one hazard having the potential to cause one or more harm(s) in the context of a domain-specific function and component. |
| Harm (ISO 14971) | Injury or damage to the health of people, or damage to property or the environment |
| Hazard (ISO 14971) | Potential source of harm. |
| Hazardous situation (ISO 14971) | Circumstance in which people, property or the environment is/are exposed to one or more hazards. |
| Risk level | Combination of probability and severity. |
| Severity (ISO 14971) | Measure of the possible consequences of a hazard. |
The Riskman ontology formalises the recent VDE Spec 90025 [7], which proposes a structured format for digitalising risk management files, as well as a machine-readable exchange format using HTML and RDFa [2]. VDE Spec 90025, in turn, is based on ISO 14971 [40], which “specifies terminology, principles and a process for risk management of medical devices”. In this section, we briefly explain the notions from VDE Spec 90025 and ISO 14971 that are central to the Riskman ontology. Some of the notions are provided in Table 1. In the supplementary material we give a full list, as VDE Spec 90025 [7] is still in press.
The model of risk underlying both ISO 14971 and VDE Spec 90025 (see Annex C of ISO 14971) is that of a hazard leading, via a sequence of events, to a hazardous situation, which, in turn, results in a harm, as depicted below:
The risk analysis that manufacturers of medical devices must undertake involves, in this view, compiling a list of known and foreseeable combinations of hazard, events, hazardous situation, and harm. A further central aspect of risk assessment is estimating (initial) risk levels by associating a probability of occurrence and severity to harms. The probability is often split into probabilities P1 and P2, with the overall probability then being .
Building on ISO 14971, VDE Spec 90025 requires manufacturers to document the results of their risk analysis as a list of analysed risks. This includes providing a device context to each analysed risk. Moreover, hazards are embedded within domain-specific hazards, which also include information on the device function and device component of the device (or class of devices from one domain, e.g. X-ray devices) to which the hazard refers. VDE Spec 90025 does not provide requirements on the values used to specify probabilities or the severity; in particular, probabilities are often not probabilities in the mathematical sense. Rather, these values are interpreted as magnitudes within a scale that is used by the manufacturing company producing the risk management file (cf. Sect. 4.4).
Risk assessment by manufacturers primarily aims at risk control, i.e. minimising risks as much as possible. To document the mitigation strategies devised for the different risks, VDE Spec 90025 proposes the use of safe design arguments (SDAs). These, although building on the notion of assurance cases [73], avoid some of the complexities in their use [45, 51, 31, 32] by only requiring a considerably simplified structure. Specifically, SDAs encode structured arguments, showing that a certain risk has been mitigated, as trees. Thus, each SDA can have one or more sub-SDAs (its children), which serve to substantiate a claim made in the parent SDA. Moreover, SDAs need ultimately be based on safe design argument implementations (SDAIs), i.e. all leaf SDAs need to be SDAIs. These include not only a claim but also an implementation manifest. The latter gives detailed information on how the claim has been implemented and points to concrete evidence (e.g. additional documentation) to support this.
VDE Spec 90025 makes a further distinction between risk SDAs and assurance SDAs, with assurance SDAs referring to some state-of-the-art safety assurance – e.g. a section of a norm or standard mentioning a way of handling a risk. All other SDAs are risk SDAs, while children of assurance SDAs need also be assurance SDAs. An SDA tree is embedded within a controlled risk, which, apart from referring to the analysed risk it “controls”, also indicates a residual risk level, i.e. probability and severity of the risk remaining after risk control.
Example 1.
Insulin infusion pumps aid in regulating blood glucose levels, especially of patients with diabetes, by administering fast-acting insulin via a catheter inserted beneath the skin. Based on the risk assessment for a generic infusion pump by Zhang et al. [75, 74], Fig. 1 shows a controlled risk and associated SDA that can be extracted from a risk management file that follows VDE Spec 90025.
forked edges,
for tree=align=center,base=bottom,draw,font=,edge=-¿,rounded corners,scale=0.9,
im/.style=fill=green!10,
sa/.style=fill=purple!15,
sda/.style=fill=blue!10,
asda/.style=fill=brown!15,
[SDA:
Alternative alerting when vibration mechanism of non-audio alarm fails (), sda,
[SDA:Additional
visual (blinking)
signal (),sda,name=sda1
[IM: Sec. 10.3
of Alarm
report (),im,name=im1]],
[SDA: Notification
recurs every
minutes (),sda,name=sda2
[IM: ,
Sec. 10.7
of Alarm
report (),im]],
[SDA: Additio-
nal audio
alarm (),sda,name=asda3
[SDA: Audio
signal if vibration
signal not
acknowledged (),sda,name=asda4
[IM: Sec.
10.11 of Alarm
report (),im]],
[SDA:
Audible signal is
at least X db
at Y m (),asda,name=asda5
[IM: ,
, Sec. 5.3
of Loudspeaker
test (),im,name=im5]],
[SA: IEC
60601 (), no edge, sa,name=sa5,yshift=6pt]
[…]
4 The Riskman Ontology & Shapes: Overview and Usage
The Riskman ontology was developed within a consortium of domain experts in medical devices and risk management, ontology engineers and software developers. For the development of the ontology, we utilised the Linked Open Terms (LOT) methodology [58], a framework for ontology development that builds upon the NeOn methodology [65]. It comprised the following activities:
Requirement specification. We studied relevant norms/standards and inspected risk management files of real medical devices to gain a shared understanding of the required terms. This was accompanied by several meetings within the consortium and workshops with domain experts from the field: risk managers, manufacturers, consultants, and notified bodies. As a result, objectives and a glossary of terms were compiled, forming the basis for the conceptual ontology model. Moreover, the VDE Spec 90025 [7] working group was formed to develop a specific submission format and its requirements based on the conceptual model.
Implementation. We encoded the ontology and shapes using OWL and SHACL. During development, we used the HermiT reasoner [63] and the OntOlogy Pitfalls Scanner (OOPS!) [59] to ensure the ontology’s validity and consistency.
Publication and Maintenance. The Riskman ontology, shapes and documentation are published online with a permanent URL provided by w3id. The ontology is maintained in a GitHub repository [8] for issue tracking and version control.
As mentioned before, the approach assumes risk documentation to be submitted as HTML files with data encoded in RDFa.
It is then extracted using an RDF distiller and used as input to an materialisation reasoner alongside with the Riskman- and potentially additional ontologies.
As a result,
an enhanced DL knowledge base is conceived, which is later validated for conformity against the Riskman SHACL constraints (and potentially additional constraints) utilising a SHACL validator.
The outcome is then communicated with a human-readable validation report.
We provide a prototypical implementation of the validation pipeline together with examples showcasing validation results [30].
The symbolical depiction below presents the architecture and data flow of the approach:
In the following subsections, we will explain the approach in more detail and provide preliminaries necessary for understanding. Afterwards, the actual Riskman ontology and shapes will be shown. This section will conclude with showcasing the approach on our running GIIP example and giving hints of how the approach can be extended depending on further needs of interested parties.
4.1 The Description Logic
Firstly, it is important to note that all of the Riskman ontology can be represented within the description logic [10, 11]. We therefore briefly recall its syntax, semantics, and aspects of reasoning that are relevant to our approach.
’s concept constructors and their semantics are recalled in the upper part of Table 2; the middle (lower) part shows the constructs allowed in a TBox (ABox).
| Name | Syntax | Semantics |
| individual name | ||
| concept name | ||
| role name | ||
| top | ||
| bottom | ||
| nominal | ||
| conjunction | ||
| existential restriction | ||
| range restriction | ||
| general concept inclusion | ||
| role inclusion axiom | ||
| concept assertion | ||
| role assertion |
As usual for description logics, the semantics of is defined via interpretations with a non-empty domain and an interpretation function . An interpretation is a model of an ABox (TBox ) iff it satisfies all elements of () as per Table 2. An assertion is entailed by , written , iff every model of is a model of . We also remark that RIAs (13)–(16) satisfy the syntactic restriction imposed by Baader et al. [11, Section 3].
There are two kinds of reasoners, predominantly ontology classifiers [43]. The first kind implements tableau calculi while the second is based on materialisation. Tools like ELK [43], also distributed with ontology modelling frontends like Protégé [53], follow the materialisation-based approach by step-wise computing all logical consequences from a given ABox (i.e., the risk management information) and an ontology (i.e., Riskman). The polynomial-time complexity of reasoning [10, 11] opens the stage for general purpose Datalog reasoners (e.g. nemo [41]) to implement materialisation [15]. Thus, a wide variety of highly optimised tools are at the disposal of potential users of Riskman.
4.2 Classes and Properties
Riskman’s classes and their interrelationships are depicted in the schema diagram in Fig. 2 (top), where also domains and ranges of properties can be read off the edges. The main axioms of the Riskman ontology are given in Fig. 2 (bottom), including detailed superclass declarations, and relationships between properties via role inclusion axioms, role chains, or transitivity statements.
4.3 Ontology Design Patterns
We opted for a lightweight ontology that captures the outcomes of performed risk management with low ontological commitment, as the particular needs might differ from manufacturer to manufacturer. To ensure the usability and extensibility of the ontology, we employed the Stub Metapattern [48, 49], which “acts as a type of placeholder for future extensions.” Specifically, the class DeviceProblem is intended to link to the IMDRF’s controlled vocabularies [39, Annex A]; another stub, HazardousSituation, enables to potentially reuse the Hazardous Situation Pattern [50, 16] should a more fine-grained representation be required.
4.4 Probabilities and Severities
VDE Spec 90025 (importing from ISO 14971) defines risk as the “combination of the probability of occurrence of harm and the severity of that harm”, and so to represent concrete risks it is necessary to also represent concrete “values” for probability (and severity). Since mathematical probabilities are virtually impossible to obtain for events and situations that are hypothetical from the outset, the typical approach in risk management (cf. Sect. 3) is to use a fixed, finite number of probability magnitudes, each representing an interval of real-valued probabilities, and being naturally ordered on a logarithmic scale, e.g. “improbable”: , “remote”: , “occasional”: , “probable”: , and “frequent”: . The exact number of different magnitudes varies and is up to the manufacturer, but choosing (as above) is common; a similar approach is typically also used to represent severity.
We refrained from binding users of the Riskman ontology to a specific way of representing probability and severity, but at the same time want to provide a reasonable baseline that can be used almost “out of the box”. To this end, we have an optional “plugin” that creates probability and severity magnitudes (ontologically represented by individuals using nominals) for given desired interval counts (probability) and (severity), together with additional axioms as an ontology (for probability-severity ontology), with where
This introduces not only the discretised probability values (e.g. for we would get that and ), but also an ordering on these values. Most importantly, however, the GCIs introduced above implement the “computation” of overall probability from probabilities and [40]. Essentially, the multiplication works by adding exponents of upper bounds of intervals, e.g. results in , and all possible computations for the given (and ) are expressed via GCIs.
4.5 Shapes
As explained above, our ontology can be used to derive implicit information via reasoning. The materialised graph is then stored in RDF format [72]. (As an illustration, we depict the implementation of Example 1 in Fig. 4.) For our use case, the evaluation of risk reports, it is now important that we are also able to check for missing information (risks could be not mitigated) or for mismatches between values (a mitigation should not increase severity or probability of a risk). These kinds of checks on a (hopefully) complete and self-contained risk report are conceptually not a good fit with OWL’s open-world assumption. We thus define SHACL shapes that operate on the materialised RDF graph and implement the most important requirements risk reports should fulfil. We describe them in what follows. Users clearly can, and most likely will, add custom shapes.
We adopt the abstract syntax of SHACL constraints proposed by Corman et al. [18, 6], capturing the core components of the SHACL specification [71]. For our purposes, we conveniently re-use description logic vocabulary, viz., pairwise disjoint sets of classes, of properties, and of individuals. A finite set of assertions (an ABox) can then be seen as representing a labelled graph, with individuals acting as nodes, classes labelling nodes, and properties labelling edges. The syntax of shape expressions and path expressions is shown in Fig. 3 (top). For the semantics, a given graph (ABox) with nodes (individuals) defines an evaluation function that assigns to each path expression a binary relation , and to each shape expression a set via structural induction as shown in Fig. 3 (middle).
A shape constraint is an expression of the form , with and a shape expression. A shape schema is a pair where is a set of shape constraints and is a set of target concept assertions. Intuitively, a target expresses the requirement that be labelled by . Formally, an ABox is a model for a set of constraints iff for all . An ABox is validated against a schema iff there exists a set of concept assertions such that (1) , (2) , and (3) is a model for .
The shapes constraints that complement the Riskman ontology are listed in Fig. 3 (bottom). To illustrate their use, in Fig. 4 individuals satisfying the shape expression part of a constraint are labelled by the constraint’s number.
The syntax of path expressions and shape expressions is given by the grammars
where , , , and with indicating the inverse of .
4.6 Example and Extensibility
The example from Fig. 4 illustrates how our ontology and shapes could be used in practice. Alas, even a carefully designed framework like ours, which takes the needs of the different stakeholders into account, can never meet everyone’s wishes. Companies could have individual requirements, authorities might want to add specialised tests, and regulations differ per region and change over time. This is one of the main reasons why we chose Semantic Web technologies for our task: they are easy to extend and can be combined with existing data. To illustrate the extensibility of our approach, we give a small example: We described earlier (Sect. 4.4) that companies often define magnitude levels to model probabilities and severities of risks. While studying risk reports, we observed that these definitions sometimes come with a so-called risk acceptance matrix, a schema indicating which combinations of probability and severity a company considers critical. Such information can be modelled with our ontology by adding a class . Assume for the sake of example that the combination “probability and severity ” is critical. This can be indicated by an axiom
allowing to easily check for controlled risks with critical risk levels using SHACL:
Of course, more complex additions and constraints are also possible.
5 Discussion and Outlook
We presented the Riskman ontology & shapes with their intended use of representing and analysing risk management information for medical devices. Analysed risks and their mitigations are represented as an ABox, the Riskman ontology is used with a materialisation engine to infer implicit knowledge, and lastly SHACL constraints are used to check whether the input data conforms to given requirements. The resource (ontology + shapes constraints) is available, including a reference implementation of the whole pipeline. With feedback from manufacturers and notified bodies incorporated into it, we envision Riskman to improve the work lives of risk managers and certification auditors alike.
The issue of medical device safety will only become more important in the future, especially with further digitisation [66]. An interesting next step for the Riskman ontology is to not only assess submission completeness, but also evaluate the quality of the assurance [17]. First steps, including a proof of concept, have recently been achieved in the context of assurance cases [28], but are known to be hard to generalise [21]. Possible techniques to consider for evaluation of risk management artifacts represented using Riskman are dialogue-based approaches to proof theory in structured argumentation [24], as well as other approaches to provide justifications in logic-based knowledge representation formalisms [23].
In the future, we also want to use existing data about adverse events (partially annotated with terms from IMDRF taxonomies, e.g. from the FDA’s Manufacturer and User Device Facility Experience database, MAUDE [70]) to allow for a more informed evaluation of “whether all hazards have been identified“, a classical point of risk management that cannot strictly be proved by a manufacturer, but potentially be disproved by a sufficiently knowledgeable reviewer/auditor.
Ahmetaj et al. [3] analysed how non-validation of SHACL constraints can be explained to users (in terms of repairs), which can potentially be applied to our work and be included in a future Riskman-based work bench for risk managers.
Instead of delegating the inference and validation steps to two different services (reasoners/validators), we could also utilise advancements in the area of combining OWL+SHACL, e.g. by converting the ontology and shapes into a single set of SHACL constraints [4]. While this involves an exponential blowup in general [4], the fact that Riskman stays within might constitute an interesting special case. Alternatively, having a single reasoner perform inferencing as well as constraint checking could also be achieved by translating ontology axioms and shapes constraints into answer set programming [14]. With different implementations in place, we will experimentally evaluate and compare them.
As a possible alternative to extending some concepts that are currently stubs (in the sense of the stub metapattern [48]), we envision to use the novel formalism of standpoint logic [33, 34] to import and attach further ontologies to Riskman. In this regard it is especially notable and useful that the combination of standpoint logic and the description logic retains the latter’s polytime computational complexity [36, 35]. A prominent candidate for integration is the US National Cancer Institute’s thesaurus (NCIt) [54], which can be expressed in [11].
Acknowledgements.
This work was supported by funding from BMBF within projects KIMEDS (grant no. GW0552B), MEDGE (grant no. 16ME0529), and SEMECO (grant no. 03ZU1210B).
Appendix 0.A Definitions of Terms from VDE Spec 90025
In this section, we briefly recall relevant definitions from VDE Spec 90025 [7]. Table 3 provides definitions of terms borrowed from ISO 14971 [40], while table 4 gives definitions of terms newly introduced in VDE Spec 90025.
| Term | Definition |
| Harm | Injury or damage to the health of people, or damage to property or the environment. |
| Hazard | Potential source of harm. |
| Hazardous situation | Circumstance in which people, property, or the environment is/are exposed to one or more hazards. |
| Intended use, intended purpose | Use for which a product, process, or service is intended according to the specifications, instructions140and information provided by the manufacturer. |
| Objective evidence | Data supporting the existence of verity of something. |
| P1 | Probability of the occurrence of a hazardous situation. |
| P2 | Probability of a hazardous situation leading to harm. |
| Residual risk | Risk remaining after risk control measures have been implemented. |
| Risk | Combination of the probability of occurrence of harm and the severity of that harm. |
| Risk analysis | Systematic use of available information to identify hazards and to estimate the risk. |
| Risk control | Process in which decisions are made and measures implemented by which risks are reduced to, or maintained within, specified levels. |
| Safety | Freedom from unacceptable risk. |
| Severity | Measure of the possible consequences of a hazard. |
| State of the art | Developed stage of technical capability at a given time as regards products, processes, and services, based on the relevant consolidated findings of science, technology, and experience. |
| Term | Definition |
| Analysed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Assurance SDA | SDA where the purpose is to make a safety assurance. |
| Assurance SDAI | SDAI of an assurance SDA. |
| Controlled risk | Structured artifact that relates one analysed risk with one or more SDA(s) and specifies a resulting residual risk. |
| Device component | A (physical or logical) part of a device. |
| Device context | Information concerning the use context of a device, including, but not limited to, (1) intended use/intended purpose, (2) instructions for use, and (3) intended environment of use. |
| Device function | Functional device capability at application level. |
| Domain-specific hazard | Structured artifact that centres around one hazard having the potential to cause one or more harm(s) in the context of a domain-specific function and component. |
| Event | Atomic occurrence or incident that (possibly when linked in a sequence with other events) may spawn a hazardous situation from a domain-specific hazard. |
| Implementation manifest | Concrete piece of objective evidence (or a reference to such) that an SDA has been implemented, e.g. reference to a line of code or a particular section in the device manual. |
| Intended environment of use | Environment or environmental conditions in which the device is intended to be used. |
| Risk matrix | Matrix (two-dimensional table) displaying all combinations of probability and severity classes without determining which of those combinations are acceptable. |
| Risk SDA | SDA where the purpose is to control a Risk. |
| Risk SDAI | SDAI of a Risk SDA. |
| Risk level | Combination of probability and severity. |
| Safety assurance | A credible reference (or list of such) to the state of the art of achieving safety with respect to a certain class of hazards, e.g. referring to an international norm such as IEC 60601. |
| SDA (Safe design argument) | Reusable artifact embodying or expressing one possible method or approach towards a specific goal. |
| SDAI (SDA implementation) | Structured artifact specifying a concrete implementation or realisation of a specific SDA. |
| Use-Context | Intended/reasonably foreseeable environment the device can be used in, that may affect a related risk. |