### v3

# Supporting Risk Management for Medical Devices via the Riskman Ontology and Shapes (Preprint)

Abstract
We propose the Riskman ontology and shapes for representing and analysing information about risk management for medical devices. Risk management is concerned with taking necessary precautions to ensure that a medical device does not cause harms for users or the environment. To date, risk management documentation is submitted to notified bodies (for certification) in the form of semi-structured natural language text. We propose to use terms from the Riskman ontology to provide a formal, logical underpinning for risk management documentation, and to use the included SHACL constraints to check whether the provided data is in accordance with the requirements of the two relevant norms, i.e. ISO 14971 and VDE Spec 90025. Riskman is available at https://w3id.org/riskman.
keywords:
[…]
Medical devices typically are safety-critical, meaning their failure under certain conditions can lead to harm to humans or the environment. To ensure that potential harms are minimized, legislation requires manufacturers of medical devices to provide a comprehensive justification that their product is acceptably safe. In Europe, such legal requirements mainly stem from the European Union’s Medical Device Regulation (EU MDR) [EUMDR]. After its introduction in 2017, the original plan was that manufacturers have time until May 2024 to re-certify their devices under the new regulations. This caused an immense backlog for notified bodies (organisations certifying medical devices in Europe, e.g., TÜVs in Germany), so the EU decided to extend the transition period until at least 2027 to avert a potential shortage of medical devices (and ramifications thereof) [mdrNewsURL].
The fundamental challenges of medical device certification are (1) the sheer amount of information that notified bodies have to process (even for a single device), and (2) the way this information is submitted to notified bodies, namely in the form of text. In principle, device manufacturers have to argue that they have proactively, systematically, and thoroughly analyzed and mitigated the risks associated with their device as much as possible. One way to do this is using a so-called assurance case (AC), a structured argument supported by a body of evidence that jointly provide a compelling, comprehensible, and valid case that a system is acceptably safe for a given application in a given context. Using assurance cases has a long(er) tradition in other areas of safety-critical systems, e.g. in aviation or nuclear power plants [SujanHKPJ16], but the need for assurance cases for medical devices has been recognized long ago [WeinstockG09]. The Food and Drug Administration of the U.S. (FDA) has even provided a guidance document for one class of products (infusion pumps) due to an unusually large number of previous incidents [FDA14].
As much as the introduction of assurance cases has helped advance safety management practices in the medical device industry, there is still a lot to be wanted. Sujan et al. even claim that “safety management practices in healthcare [are] at present […] less mature than those in other safety-critical industries” [SujanHKPJ16, p. 185]. We briefly note some identified shortcomings:
(1) Since the intended recipient of a safety case is human (the auditor), a lot of work in safety engineering has focused on the presentation of the argument, e.g. to avoid confirmation bias [ChowdhuryWPL20]. Several approaches to visualize assurance cases exist, e.g. the goal structuring notation [KellyM97]. However, work on the representation of safety arguments is rare, since applicable legislation typically requires a submission in the form of text.
(2) Due to the sheer volume of text submitted to notified bodies for certification, an auditor can never look at all relevant points in detail; furthermore, auditors typically spend a considerable proportion of their time navigating through documents (in the best case using text search in a document viewer). Tedious, manual examination of submissions is necessary to identify missing or incorrect information. Significant time savings are offered when the initial inspection can be (semi-)automated and when improved navigation through e.g. semantics-enhanced document search is available.
(3) Reusing parts of safety arguments is a widely recognized problem [MartinBLW14, RuizJEVL17]. The simplest instance of re-use happens when a manufacturer continues development of a certified device and wants to certify the “updated version”: application lifecycle management tools help identify the delta between two device versions and risk managers can restrict attention on how that difference affects risks and mitigations. Harder instances of re-use are equally pervasive and important. For example, this occurs when a manufacturer wants to use some device part in a novel device of a different type. It also arises when a manufacturer buys parts (like integrated circuits) from component suppliers and uses assurances about those parts to claim overall device safety. We even propose to go beyond this and technically enable re-use across manufacturers, which is especially relevant for small/medium enterprises that may lack staff or experience in risk management.
Therefore, any approach toward a more structured representation of safety cases, allowing for automated checks and facilitating re-use, would appear to be highly welcome by all involved parties. Consequently, in this paper, we propose to use logical modelling (more precisely the web ontology language, OWL) to represent risk management documentation, and the Shapes Constraint Language (SHACL) to check those representations for conformance with a set of requirements, e.g., whether all identified risks have an associated mitigation. While such straightforward checks could clearly also be achieved by custom software, the presence of a logical inference step before the constraint checking step is a significant advantage of our approach, as we shall demonstrate later in the paper. Thus as our main contribution we introduce the Riskman ontology and shapes.
As far as we are aware, using OWL and SHACL is a novel approach to representing and reasoning about risk management documentation. The requirements to be checked are simple and syntactic at the moment, but in view of communication with domain experts from notified bodies and manufacturers, we still expect our approach to be a major step forward as manufacturers (and notified bodies) can expect to the first round of conformity assessment being fully automated. We would like to emphasize that our approach does not aim at assessing the adequacy and correctness of the implemented mitigations themselves (such as whether an insulation thickness of 0.5 mm is adequate to lower the risk of electric shock). It stands to reason that such an analysis is presumably AI-complete and would at least require vast amounts of background knowledge. Apart from facilitating tasks in this important application domain, our ontology and shapes are also of independent interest to the Semantic Web community, as the combination of logical inference and SHACL constraint checking is a topic that has garnered significant research interest lately [AhmetajOOS23, ParetiKNS19, demeesterswj2021]. Finally, the fact that our approach relies on standardized and well-supported formalisms unlocks the adoption of Semantic Web technologies in general. Employing these technologies enables Riskman users to benefit from existing infrastructure of tooling: Risk reports can easily be queried or enhanced with additional information about the use context of a concrete device or data about particular patient problems by simply using other ontologies and data [cc, el2018snomed, ncitURL].
The rest of the paper is organized as follows: In the next section, we give an overview on related work. Afterwards (Section 3), we introduce the state of the art of risk management and the basic notions introduced by the norms upon which we build our ontology, both in general terms and with an illustrative running example. Section 4 then introduces the Riskman ontology and its associated shape constraints, explaining how we intend them to be used in risk management, and showcasing them with the running example. In Section 5 we conclude with a discussion of potential future work.
2 Related Work
We reviewed the literature for (1) ontologies on (a) medical devices or (b) notions of risk, with a special focus on an intended use for automated reasoning (in particular using SHACL constraints), and (2) academic reports on software tools for risk management.
2.1 Conceptual Work and Ontologies
Fenz et al. [FenzPH16, FenzN18] created an ontology for the information security standard ISO 27002. While they do not deal with risk management (or medical devices), they also use their ontology along with a reasoner to infer information about compliance with the standard’s requirements. Reasoning results are then interpreted by humans or by a tailor-made software tool. Uciteli et al. [UciteliNTSSFKSS17] provided the Risk Identification Ontology, which defines notions of risk and adverse situation and embeds them into the top-level ontology GFO (General Formal Ontology) [Herre10, HerreHBHLM06]. However, they work in a more process-oriented setting with special focus on risk identification in time periods surrounding surgical procedures. In that setting, risks cannot be mitigated beforehand, and so risk management – our main focus – is not within their scope. Kim et al. [KimPLL19] presented a process integration ontology for medical software developers with a focus on medical devices, combining notions from IEC 60601-1, IEC 62304, and, notably, ISO 14971. The resulting integrated ontology is however mainly designed to help developers comply with the involved standards; the ontology was not developed with logical reasoning as explicit intended use case. Aziz et al. [AzizAK19] developed a Hazard Identification Ontology, involving notions of Hazard and Events. Their focus was however on identifying risks rather than mitigating them, especially in scenarios of fire, explosions, or toxicity. Schütz et al. [SchutzFW20] created an ontology for medical devices in Germany, albeit more broadly targeting devices’ manufacturers, operators, and legal procedures from an outside perspective with the aim of general semantic interoperability, and based on a legal framework that has since been superseded by the EU MDR. Single et al. [SingleSD20] presented an ontology for Hazard and Operability (HAZOP), a methodology for scenario-based hazard evaluation, therein defining notions of deviation, cause, effect, consequence, and safeguard. Their aim was however to create HAZOP worksheets (to be used by human operators) automatically. McDonald et al. [McDonald21] mention how they use the “ARK Mindful Governance of operational risk formal ontology” to annotate textual risk-analysis data to make it amenable to machine processing (p12). They however give no details on the ontology or its possible inferences. Alanen et al. [AlanenLMPAHT22] provided a comprehensive risk assessment ontology including the notions risk, risk level, and risk control, which is harmonized between (and intended for use across) safety, security, and dependability. The intended use of their ontology is to “support the creation of a structured work product storage with traceability links” in order to improve upon current practices with non-structured word-processing documents, albeit without using logical reasoning. Golpayegani et al. [GolpayeganiPL22] propose the AI Risk Ontology (AIRO) for expressing information associated with high-risk AI systems based on the requirements of the EU’s AI Act [AIAct] and the ISO 31000 series of standards. AIRO contains classes for risk, risk source, event, consequence, and impact, and Golpayegani et al. also present SHACL shapes for determining whether an AI system is high-risk. This is perhaps the closest relative of Riskman, only for AI systems instead of medical devices. In conclusion, while various ontologies including notions of risk exist, they either do not implement ISO 14971 [ISO14971] or are not designed with logical reasoning in mind, and most of them do not use SHACL for conformance testing (in the sense of pre-certification).
2.2 (Logic-Based) Software Tools for Risk Management
Fujita et al. [FujitaMHSKI12] presented the D-Case Editor, a dependability-oriented assurance case editor supporting the goal structuring notation [KellyM97], implemented as an Eclipse plugin and available at its GitHub repository [d-case_editorURL]. However, its development has been discontinued. Rushby [Rushby05] and Cruanes et al. [CruanesHOS13] introduced the Evidential Tool Bus, a Datalog-based system for integrating the development of safety cases into system (and software) development. In particular, their approach covers the management of claims (about the system to be developed) and how they are supported by evidence provided by other software tools. While an implementation is available [ETBURL], it is not maintained any longer. Beyene and Ruess [BeyeneR18] later picked up that work and extended it by connecting to the Jenkins [JenkinsURL] continuous integration software, with an active repository [JenkinsURLGit]. The focus of the Evidential Tool Bus is primarily on obtaining, maintaining, and managing pieces of evidence in a development setting. As such, it can be an important and complementary addition to using the methodology we propose in this paper. Since the issue of safety is of great concern to industry, there are also several proprietary software tools centring around safety assurance: The UK company Adelard offers the Assurance and Safety Case Environment. Similarly, de la Vara et al. [delaVaraJMP19] reported on using the V&V studio by REUSE Software. In conclusion, existing software solutions for risk management are proprietary or do not use formal logics, or they have other (but complementary) use cases.
3 Background
| Term | Definition |
| Analyzed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Controlled risk | Structured artifact that relates one analyzed risk with one or more SDA(s) and specifies a resulting residual risk. |
| Domain specific hazard | Structured artifact that centres around one hazard having the potential to cause one or more harm(s) in the context of a domain-specific function and component. |
| Harm (ISO 14971) | Injury or damage to the health of people, or damage to property or the environment |
| Hazard (ISO 14971) | Potential source of harm. |
| Hazardous situation (ISO 14971) | Circumstance in which people, property or the environment is/are exposed to one or more hazards. |
| Risk level | Combination of probability and severity. |
| Severity (ISO 14971) | Measure of the possible consequences of a hazard. |
The Riskman ontology is based on the recent VDE Spec 90025 [VDESpec24], which proposes a structured format for digitalizing risk management files, as well as a machine-readable exchange format using HTML with RDFa [RDFa], which annotates (some) HTML tags with Resource Description Framework (RDF) triples. VDE Spec 90025, in turn, is based on ISO 14971 [ISO14971], which “specifies terminology, principles, and a process for risk management of medical devices”. Throughout this section, we briefly recall the notions from VDE Spec 90025 and ISO 14971 that are central to the Riskman ontology. Some of the notions are provided in Table 1. A full list of definitions of the terms can be found in VDE Spec 90025 [VDESpec24, Section 3]. The model of risk underlying both ISO 14971 and VDE Spec 90025 (see Annex C of ISO 14971) is that of a hazard leading, via a sequence of events, to a hazardous situation, which, in turn, results in a harm, as depicted below:
According to this view, the risk analysis that manufacturers of medical devices must undertake involves compiling a list of known and foreseeable combinations of hazard, events, hazardous situation, and harm. A further central aspect of risk assessment is estimating (initial) risk levels by associating a probability of occurrence and a severity to each harm. The probability is often split into probabilities and , with the overall probability then being . Building on ISO 14971, VDE Spec 90025 requires manufacturers to document risk analysis results as a list of analyzed risks, including related harm, patient problem, device context, the hazardous situation caused by preceding event(s), the assessed initial risk level, and the associated domain specific hazards. These, in turn, further specify the underlying hazard through details on the device problem, device function, and device component. The “problem” classes enable referencing terminology from the International Medical Device Regulators Forum (IMDRF) [IMDRF20AET], specifically Annex E for patient problems and Annex A for device problems. The notion of a device is absent in VDE Spec 90025 and the Riskman ontology, since each risk management file addresses a single medical device. However, extending the ontology to include a device concept and link related entities would enable more comprehensive queries. VDE Spec 90025 does not provide requirements on the values used to specify probabilities or the severity either; in particular, probabilities are not probabilities in the mathematical sense. Rather, these values are interpreted as magnitudes within a scale that is used by the manufacturing company producing the risk management file (cf. Section 4.3).
Risk assessment by manufacturers primarily aims at risk control, i.e. minimizing risks as much as possible [EUMDR]. To document the mitigation strategies devised for the different risks, VDE Spec 90025 proposes the use of safe design arguments (SDAs). These, although building on the notion of assurance cases [WeinstockG09], avoid some complexity in their use [Kelly03, Leveson11, Graydon15, Graydon17] by only requiring a considerably simplified structure. Specifically, SDAs encode structured arguments, showing that a certain risk has been mitigated, as trees. Thus, each SDA can have one or more sub-SDAs (its children), which serve to substantiate a claim made in the parent SDA. Moreover, SDAs need ultimately be based on safe design argument implementations (SDAIs), i.e. all leaf SDAs need to be SDAIs. These include not only a claim but also an implementation manifest, which gives detailed information on how the claim has been implemented and points to concrete evidence (e.g. additional documentation or a specific line of software source code) to support this.
VDE Spec 90025 further distinguishes between risk SDAs and assurance SDAs, with assurance SDAs referring to some state-of-the-art safety assurance – e.g. a section of a norm mentioning a way of handling a risk. All other SDAs are risk SDAs, while children of assurance SDAs must also be assurance SDAs. An SDA tree is embedded within a controlled risk, which, apart from referring to the analyzed risk it “controls”, also indicates a residual risk level, i.e. probability and severity that remain after risk control.
Example 1.
Insulin infusion pumps aid in regulating blood glucose levels, especially of patients with diabetes, by administering fast-acting insulin via a catheter inserted beneath the skin. Based on the risk assessment for a generic infusion pump by Zhang et al. [zhangJJ10, zhangJJR11], Figure 1 shows a controlled risk and associated SDA that can be extracted from a risk management file that follows VDE Spec 90025. Following Zhang et al. [zhangJJ10, entry 4.3.9 in Table 4 in the appendix], the risk stems from an “alarm malfunction” hazard (indicated by in Figure 1). It is further specified within a domain specific hazard () with related information on the device component (), its function () and problem (). Specifically, the vibration mechanism of the non-audio alarm integrated into the pump may fail (event ) under normal operating conditions (device context ). Then, the patient may not become aware of an issue (event ), which can lead to the patient receiving less insulin (hazardous situation ) and the patient losing consciousness (harm ). The SDA (, based on the work of Zhang et al. [zhangJJR11, Table 3]) consists of three sub-SDAs and expresses that there are alternative means of alerting the patient. The first sub-SDA () expresses that the alarm condition is also indicated through visual signals; the second sub-SDA () indicates that this notification is recurring; the third sub-SDA () expresses that there is also an additional audio alarm that will start unless the patient acknowledges the vibration or blinking. Moreover, according to sub-SDA () the audible signal is in accordance with regulations, here the assurance is IEC 60601 (). Thus, sub-SDA () is the only assurance SDA; all other SDAs are risk SDAs. On the other hand, as required by VDE Spec, each leaf SDA is an SDAI, with their implementation manifests (,,,) pointing to details and documentation.
4 The Riskman Ontology and Shapes: Overview and Usage
The Riskman ontology was developed taking into account requirements and conceptual design choices laid out by a consortium of domain experts in medical devices and risk management, ontology engineers, and software developers. For ontology development, we used the Linked Open Terms (LOT) methodology [Poveda-Villalon2022], an ontology development framework that builds upon the NeOn methodology [SurezFigueroa2015TheNM]. It comprised the following:
Requirement specification. Studying relevant norms/standards and inspecting real medical devices’ risk management files to gain a shared understanding of the required terms. This was accompanied by several meetings within the consortium and workshops with domain experts from the field: risk managers, manufacturers, consultants, and notified bodies. As a result, we compiled objectives and a glossary of terms, forming the basis for the conceptual ontology model. In parallel, the VDE Spec 90025 [VDESpec24] working group developed a technical submission format with constraints based on the conceptual model.
Implementation. We encoded the ontology and shapes using OWL and SHACL. During development, we used the HermiT reasoner [shearer2008hermit] and the OntOlogy Pitfalls Scanner (OOPS!) [poveda2014oops] to ensure the ontology’s validity and consistency.
Publication and Maintenance. The Riskman ontology, shapes, and documentation are published online with a permanent URL provided by w3id. To assess the conformity of the ontology to the FAIR (Findable, Accessible, Interoperable, and Reusable) principles [wilkonson2016fair], we used the Ontology Pitfall Scanner for the FAIR Principles (FOOPS!) [foops2021]. Riskman is maintained in a GitHub repository [RiskmanRepo] for issue tracking/version control.
As previously noted, the approach expects risk documentation to be submitted as HTML files containing risk management data encoded as RDF triples.
An RDF distiller extracts the RDF graph from the HTML by removing the markup and isolating the RDF triples, which are then used as input to an reasoner alongside Riskman (and potentially additional ontologies).
As a result,
a materialized DL knowledge base is obtained, which is then validated against the Riskman SHACL constraints (and potentially additional constraints) utilizing a SHACL validator.
The outcome is then communicated with a human-readable validation report.
We provide a prototypical implementation of the validation pipeline together with examples showcasing validation results [RiskmanPipeline].
The symbolical depiction below presents the architecture and data flow of the approach:
In the following subsections, we explain the approach in more detail and provide preliminaries necessary for understanding. Afterwards, we present the actual Riskman ontology and shapes. The section concludes with illustrating the approach on our running example and showing how it can be extended to further needs of interested parties.
4.1 The Description Logic
Firstly, it is important to note that all the Riskman ontology can be represented within the description logic [BaaderBL05, BaaderLB08]. We therefore briefly recall its syntax, semantics, and aspects of reasoning that are relevant to our approach.
’s concept constructors and their semantics are recalled in the upper part of Table 2; the middle (lower) part shows the constructs allowed in a TBox (ABox).
| Name | Syntax | Semantics |
| individual name | ||
| concept name | ||
| role name | ||
| top | ||
| bottom | ||
| nominal | ||
| conjunction | ||
| existential restriction | ||
| range restriction | ||
| general concept inclusion | ||
| role inclusion axiom | ||
| concept assertion | ||
| role assertion |
As usual for description logics, the semantics of is defined via interpretations with a non-empty domain and an interpretation function . An interpretation is a model of an ABox (TBox ) if it satisfies all elements of () as per Table 2. An assertion is entailed by , written , if every model of is a model of . We also remark that RIAs (12)–(15) satisfy the syntactic restriction imposed by Baader et al. [BaaderLB08, Section 3].
There are two kinds of reasoners, predominantly ontology classifiers [KKS:ELK2013]. The first kind implements tableau calculi while the second is based on materialisation. Tools like ELK [KKS:ELK2013], also distributed with ontology modelling frontends like Protégé [Musen15], follow the materialisation-based approach by step-wise computing the relevant logical consequences of a given ABox and an ontology. In the context of Riskman, the “relevant” consequences are all concept and role assertions involving concept, role, and individual names occurring in the input ABox, where concept assertions must be atomic, that is, of the form with . The fact that no new individuals are required for this process is guaranteed by Riskman’s design, which employs existential quantification only on the left-hand side of general concept inclusion axioms.111The optional add-on of Section 4.3 has existential restriction on right-hand sides, but only involving nominal concepts, thus no new individuals need to be created during inference. The completion (addition of all relevant consequences) of a given ABox w.r.t. the Riskman ontology can therefore (also due to the satisfaction of the syntactic restriction by Baader et al. [BaaderLB08, Section 3]) be computed in polynomial time [BaaderBL05, BaaderLB08], which enables general purpose Datalog reasoners (e.g. Nemo [nemo24]) to implement Riskman ABox completion [CDK20:DatalogEL]. Thus overall, a variety of optimized tools are at the disposal of potential Riskman users.
4.2 Classes, Properties, and Ontology Design Patterns
The main axioms of the Riskman ontology are given in Figure 2, including detailed superclass declarations, and relationships between properties via role inclusion axioms. Riskman’s classes and their general interrelationships are depicted in the schema diagram in Figure 3, where also domains and ranges of properties can be read off the edges.
We opted for a lightweight ontology that captures the outcomes of risk management with low ontological commitment, as the particular needs might differ from one manufacturer to another. To ensure the usability and extensibility of the ontology, we employed the Stub Metapattern [KrisnadhiH16, StubURL], which “acts as a type of placeholder for future extensions.” Specifically, the class DeviceProblem is intended to link to the IMDRF’s controlled vocabularies [IMDRF20AET, Annex A]; another stub, HazardousSituation, enables to reuse the Hazardous Situation Pattern [Lawrynowicz2015TheHS, Cheatham2016AMT] if a more fine-grained representation is required.
4.3 Probabilities and Severities
VDE Spec 90025 (importing from ISO 14971) defines risk as the “combination of the probability of occurrence of harm and the severity of that harm”, and so to represent concrete risks it is necessary to also represent concrete “values” for probability (and severity). Mathematical probabilities are virtually impossible to accurately determine for events and situations that are hypothetical from the outset. Therefore, the typical approach in risk management (cf. Section 3) is to use a fixed, finite number of probability magnitudes, each representing an interval of real-valued probabilities, and being naturally ordered on a logarithmic scale, e.g. “improbable”: , “remote”: , “occasional”: , “probable”: , and “frequent”: . The exact number of different magnitudes varies and is up to the manufacturer; choosing (as above) is common. A similar approach is typically also used for severity.
We refrained from binding users of the Riskman ontology to a specific way of representing probability and severity, but at the same time want to provide a reasonable baseline that can be used almost “out of the box”. To this end, we have an optional “plugin” that creates probability and severity magnitudes (ontologically represented by individuals using nominals) for given desired interval counts (probability) and (severity), together with additional axioms as an ontology (for probability-severity ontology), with where
This introduces not only the discretized probability values (e.g. for we get and ), but also an ordering on these values. Most importantly, however, the GCIs in implement the “computation” of overall probability from probabilities and [ISO14971]. Essentially, the multiplication works by adding exponents of upper bounds of intervals, e.g. yields , and all possible computations for the given are expressed via GCIs.
4.4 Shapes
As explained above, our ontology can be used to derive implicit information via reasoning. The materialized graph is then stored in RDF format [owlov]. (As an illustration, we depict the implementation of 1 in Figure 5.) Our use case, the evaluation of risk reports, requires the capability to check for missing information or for mismatches between values (e.g., whether a mitigation of a risk does not increase its severity or probability). These kinds of checks on a (hopefully) complete and self-contained risk report are conceptually not a good fit with OWL’s open-world assumption. We thus define SHACL shapes that operate on the materialized RDF graph and implement the most important requirements risk reports should fulfil. We describe them in what follows.
We adopt the abstract syntax of SHACL constraints proposed by Corman et al. [CormanRS18, AndreselCORSS20], capturing the core components of the SHACL specification [SHACL]. For our purposes, we conveniently re-use description logic vocabulary, viz., pairwise disjoint sets of classes, of properties, and of individuals. A finite set of assertions (an ABox) can then be seen as representing a labelled graph, with individuals acting as nodes, classes labelling nodes, and properties labelling edges. The syntax of shape expressions and path expressions is shown in Table 3 (top). For the semantics, a given graph (ABox) with nodes (individuals) defines an evaluation function that assigns to each path expression a binary relation , and to each shape expression a set via induction as shown in Table 3 (bottom).
The syntax of path expressions and shape expressions is given by the grammars
where , , , and with indicating the inverse of .
A shape constraint is an expression of the form , with and a shape expression. A shape schema is a pair where is a set of shape constraints and is a set of target concept assertions. Intuitively, a target expresses the requirement that be labelled by . Formally, an ABox is a model for a set of constraints if for all . An ABox is validated against a schema if there exists a set of concept assertions such that (1) , (2) , and (3) is a model for . The shapes constraints that complement the Riskman ontology are listed in Figure 4. To illustrate their use, in Figure 5 individuals satisfying the shape expression part of a constraint are labelled by the constraint’s number.
Constraint 19 encodes checks for non-increasing residual risk levels, i.e. whether the probability or severity after implementing a mitigation is not higher than before. Constraints 16, 18, 20, and 21 are duals of the subclass declarations for , , , and , respectively, in that they enforce instances of these classes to contain all the necessary components of the class definitions, as required by VDE Spec 90025. Constraint 17 encodes that (i) every sub- of an must be an and (ii) an must have a . Assuming the ABox has previously been materialized not only by means of the Riskman ontology, but also the additional “probability-severity” ontology , Constraint 21 requires that (i) an overall probability is present (specified directly or inferred from probabilities and via ), (ii) that holds in case all three have been specified, and (iii) a severity magnitude is present. Finally, the check whether all leaf nodes of the SDA tree are s is provided by Constraint 22. Figure 5 illustrates how to use our ontology and shapes by implementing the example of Figure 1.
4.5 Extensibility
Alas, even a carefully designed framework like ours, which takes the needs of the different stakeholders into account, can never meet everyone’s wishes. Companies could have individual requirements, authorities might want to add specialized tests, and regulations differ per region and change over time. To illustrate the extensibility of our approach, we give an example: We described in Section 4.3 that companies often define magnitude levels to model probabilities and severities of risks. In practice, these definitions sometimes come with a so-called risk acceptance matrix, a schema indicating which combinations of probability and severity a company considers critical. This can be modelled with our ontology by adding a class . Assume now for the sake of example that combination “probability and severity ” is critical. This can be stated via an axiom allowing to easily check for controlled risks with critical residual risk levels using SHACL via a constraint . Of course, more complex additions of further shapes and constraints are likewise possible.
5 Discussion and Outlook
We presented the Riskman ontology and shapes with their intended use of representing and analysing risk management information for medical devices. Analyzed risks and their mitigations are represented as an ABox, the Riskman ontology is used with a reasoner to infer implicit knowledge, and lastly SHACL constraints are used to check whether the input data conform to given requirements. The ontology and shapes constraints are freely available, including a reference implementation of the whole pipeline. With feedback from manufacturers and notified bodies incorporated into it, we envision Riskman to improve the work lives of risk managers and certification auditors alike.
The issue of medical device safety will only become more important in the future, especially with further digitization [SujanSC19]. This is the case even more so for devices that use artificial intelligence themselves, and as such fall under the regulations of the AI Act [AIAct]; an integration with the like-minded AIRO ontology [GolpayeganiPL22] is an important topic for future work. An interesting next step for the Riskman ontology is to not only assess submission completeness, but also evaluate the quality of the assurance [ChowdhuryWPL20]. First steps, including a proof of concept, have recently been achieved in the context of assurance cases [FosterNGWK21], but are known to be hard to generalize [delaVaraJMP19]. Possible techniques to consider for evaluation of risk management artifacts represented using Riskman are dialogue-based approaches to proof theory in structured argumentation [DillerGG21], as well as other approaches to provide justifications in logic-based knowledge representation formalisms [DeneckerBS15].
Ahmetaj et al. [AhmetajDOPSS21] analyzed how non-validation of SHACL constraints can be explained to users (in terms of repairs), which can potentially be applied to our work and be included in a future Riskman-based work bench for risk managers.
Instead of delegating the inference and validation steps to two different services (reasoners/validators), we could also utilize advancements in the area of combining OWL and SHACL, e.g. by converting the ontology and shapes into a single set of SHACL constraints [AhmetajOOS23]. While this involves an exponential blowup in general [AhmetajOOS23], the fact that Riskman stays within might constitute an interesting special case. Alternatively, having a single reasoner perform inferencing as well as constraint checking could also be achieved by translating ontology axioms and shapes constraints into answer set programming [BrewkaET11]. Different implementations could then be compared experimentally.
As a possible alternative to extending some concepts that are currently stubs (in the sense of the stub metapattern [KrisnadhiH16]), we envision to use the novel formalism of standpoint logic [AlvarezR21, AlvarezRS22] to import and attach further ontologies to Riskman. In this regard it is especially notable and useful that the combination of standpoint logic and the description logic retains the latter’s polynomial time computational complexity [AlvarezRS23a, AlvarezRS23b]. A prominent candidate for integration is the US National Cancer Institute’s thesaurus (NCIt) [ncitURL], which can be expressed in [BaaderLB08]. On the other hand, Riskman itself could also be embedded into a top-level ontology, e.g. BFO [Arp15] or GFO [Herre10].
Acknowledgements.
This work was supported by funding from BMFTR (Federal Ministry of Research, Technology and Space) within projects KIMEDS (grant no. GW0552B), MEDGE (grant no. 16ME0529), SEMECO (grant no. 03ZU1210B), and SECAI (within DAAD project 57616814, School of Embedded Composite AI, https://secai.org/, as part of the program Konrad Zuse Schools of Excellence in Artificial Intelligence).
We are indebted to numerous people who explained the theory and practice of risk management, commented on our earlier formalisations, and participated in discussions that improved our understanding of the domain. In alphabetical order, we thank: Philipp Bank, Manuel Baur, Samet Bayraktar, Gustav Bieberstein, Nidhal Chouchane, Tony Dietrich, Anne Eßlinger, Ludger Evers, Felix Gebhardt, Stephen Gilbert, Andreas Halbleib, Christian Helmbold, Evi Hartig, Aleksandr Ilinykh, Abtin Jamshidirad, Simon Kilcher, Jürgen Koch, Tina Küttner, Andreas Lämmerzahl, Felix Lempke, Alf Ludwig, Bokhodir Mamadaliev, Svetlana Miasoedova, Davood Moghadas, Martin Neumann, Andreas Purde, Katharina Rehde, Christian Rosenzweig, Kerstin Rothweiler, Jasmine Schirmer, Hubert Sefkovicz, Sebastian Schostek, Robin Seidel, Robert Stelzmann, Sarah Tsurkan, Heike Vocke, Hans Wenner, Martin Witte, Sven Wittorf, Juliane Wober, and Uwe Zeller.
Appendix A Definitions of Terms from VDE Spec 90025
In this section, we briefly recall relevant definitions from VDE Spec 90025 [VDESpec24]. Table 4 provides definitions of terms borrowed from ISO 14971 [ISO14971], while table 5 gives definitions of terms newly introduced in VDE Spec 90025.
| Term | Definition |
| Harm | Injury or damage to the health of people, or damage to property or the environment. |
| Hazard | Potential source of harm. |
| Hazardous situation | Circumstance in which people, property, or the environment is/are exposed to one or more hazards. |
| Intended use, intended purpose | Use for which a product, process, or service is intended according to the specifications, instructions140and information provided by the manufacturer. |
| Objective evidence | Data supporting the existence of verity of something. |
| Probability of the occurrence of a hazardous situation. | |
| Probability of a hazardous situation leading to harm. | |
| Residual risk | Risk remaining after risk control measures have been implemented. |
| Risk | Combination of the probability of occurrence of harm and the severity of that harm. |
| Risk analysis | Systematic use of available information to identify hazards and to estimate the risk. |
| Risk control | Process in which decisions are made and measures implemented by which risks are reduced to, or maintained within, specified levels. |
| Safety | Freedom from unacceptable risk. |
| Severity | Measure of the possible consequences of a hazard. |
| State of the art | Developed stage of technical capability at a given time as regards products, processes, and services, based on the relevant consolidated findings of science, technology, and experience. |
| Term | Definition |
| Analyzed risk | Combination of one or more domain-specific hazard(s) with one hazardous situation and one harm with reference to a device context and a specification of an initial risk level. |
| Assurance SDA | SDA where the purpose is to make a safety assurance. |
| Assurance SDAI | SDAI of an assurance SDA. |
| Controlled risk | Structured artifact that relates one analyzed risk with one or more SDA(s) and specifies a resulting residual risk. |
| Device component | A (physical or logical) part of a device. |
| Device context | Information concerning the use context of a device, including, but not limited to, (1) intended use/intended purpose, (2) instructions for use, and (3) intended environment of use. |
| Device function | Functional device capability at application level. |
| Domain-specific hazard | Structured artifact that centres around one hazard having the potential to cause one or more harm(s) in the context of a domain-specific function and component. |
| Event | Atomic occurrence or incident that (possibly when linked in a sequence with other events) may spawn a hazardous situation from a domain-specific hazard. |
| Implementation manifest | Concrete piece of objective evidence (or a reference to such) that an SDA has been implemented, e.g. reference to a line of code or a particular section in the device manual. |
| Intended environment of use | Environment or environmental conditions in which the device is intended to be used. |
| Risk matrix | Matrix (two-dimensional table) displaying all combinations of probability and severity classes without determining which of those combinations are acceptable. |
| Risk SDA | SDA where the purpose is to control a Risk. |
| Risk SDAI | SDAI of a Risk SDA. |
| Risk level | Combination of probability and severity. |
| Safety assurance | A credible reference (or list of such) to the state of the art of achieving safety with respect to a certain class of hazards, e.g. referring to an international norm such as IEC 60601. |
| SDA (Safe design argument) | Reusable artifact embodying or expressing one possible method or approach towards a specific goal. |
| SDAI (SDA implementation) | Structured artifact specifying a concrete implementation or realisation of a specific SDA. |
| Use-Context | Intended/reasonably foreseeable environment the device can be used in, that may affect a related risk. |